# Testing Bell inequalities and probing quantum entanglement at a muon collider

### v1

@@ III   Numerical Simulation and results
We examine the sensitivity of a future Muon Collider to the entanglement and violation of Bell inequality in $H\to ZZ$ process. The corresponding Higgs boson is produced through VBF process: $W^{+}W^{-}\rightarrow H$. It is worth noting that the signal of the VBF process in a Muon Collider is very clean, almost background free as to be shown below. Here the Higgs boson is on-shell and eventually decays into pair of $Z$ bosons of which one is on-shell and the other being off-shell. We perform a Monte-Carlo simulation of $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}H$, $H\to ZZ^{*}\to(\ell^{+}\ell^{-})(\ell^{+}\ell^{-})$ events, where $\ell\in\{e,\mu\}$ and $Z^{*}$ representing the off-shell $Z$ boson, using the Madgraph software [40] which includes full spin correlation and Breit-Wigner effects. A sample of one million events are generated at leading order with $\sqrt{s}=1,3,10$ TeV energies for a Muon Collider, then the events are passed to PYTHIA 8.3 [48] for further event hadronization. Eventually all events are passed through Delphes program [41] to include detector effects with a configured file designed for a Muon Collider  [^1] https://github.com/delphes/delphes/blob/master/cards/delphes_card_MuonColliderDet.tcl.
@@
The LO cross sections for higgs production through VBF is relatively large at a TeV Muon Collider [50], which is about $1{\rm pb}$. However, the final cross sections for the considered process is suppressed due to the smaller branching ratio of the Higgs boson into four leptons (muons or electrons). The final leptons states coming from the $Z$ boson pairs can be identified as four electrons, four muons or two electrons and two muons. We focus on the $e^{+}e^{-}\mu^{+}\mu-$ category for simplicity. We choose not to apply cuts on final state leptons to avoid possible bias in quantum entanglement measurement, when generating parton-level events with Madgraph . On the other hand, cuts on the kinematic variables such as transverse momentum $p_{\rm T}$ and pseudo-rapidity $\eta_{\ell}$ of leptons are applied when running Delphes program to mimic the real experiment detector effect. The minimum value of the electron or muon transverse momentum, $p_{\rm T}$, is set to 0.5 GeV, and the maximum value for absolute rapidity, $|\eta_{\ell}|$, is 2.5.
@@
In papers [10, 27, 42], it is argued that the two $Z$ boson states is not differentiable because of their interaction nature with leptons. However, this difficulty may be overcame in the process we are working with. The on-shell $Z$ boson can be identified if the invariant mass of lepton pairs is very close to the true physical mass of $Z$ boson. Similarly, the off-shell $Z$ can also be identified if the lepton invariant mass is much smaller than actual $Z$ mass. This may help us differentiate decay products coming $Z$ boson pairs. For convenience, we label on-shell $Z$ as $Z_{1}$ whose momentum are determined from summing over its decay products $\ell_{1}^{+}\ell_{1}^{-}$, while labelling the off-shell $Z$ as $Z_{2}$ and its momentum are reconstructed from its decay leptons $\ell_{2}^{+}\ell_{2}^{-}$. We apply identical coordinate system to measure the spin polarization observable as in references [4, 25]. First of all, a set of coordinates are set up in Higgs boson rest frame from which another set of coordinate in $Z$ boson rest frame is obtained via a rotation. Then a Lorentz boost is applied on the momentum of final leptons so that they are now in the Z boson rest frame. Finally, the angular coordinates of the leptons, $(\theta_{1}^{-},\phi_{1}^{-})$ for negative-charged lepton from $Z_{1}$ decay and $(\theta_{2}^{-},\phi_{2}^{-})$ for negative-charged lepton from $Z_{2}$ decay can be obtained and used to determine the coefficients for density matrix in Eq. 12.
@@
Figure 3: The expectation value of Bell inequality as function of the off-shell Z mass $M_{Z}^{*}$. $I_{3}$ is measured at three CM energy. Points represent mean value of $I_{3}$.
Backgrounds at muon colliders are generally several orders of magnitude smaller than at hadron colliders [51]. We expect the electroweak process $\mu^{+}\mu^{-}\to ZZ\to 4\ell$ to be the main background in our analysis, where the two Z bosons are produced through W boson fusion. The 4 lepton invariant mass distributions for the signal and background are shown in Fig. 2. The background events are clearly separated from the signal events, and so it is a very good approximation to assume that there is no background to our Higgs signal.
The statistical uncertainties of the coefficients $C_{2,1,2,-1},~{}C_{2,2,2,-2}$ and $I_{3}$ are estimated through pseudo-experiments based on the event yields are determined for the target luminosity. The central values are calculated from the differential distributions, which functions of spherical coordinates of leptons. We can obtain the mean and standard deviation through repeating over pseudo-experiments. Here, the standard deviation is understand as statistical uncertainty. Because of the clean final states and fine experimental resolutions for light charged leptons, systematic uncertainty are not considered in this study. As stated in previous sections, the entanglement strength is relevant with the off-shell Z boson mass, the larger $M_{Z}^{*}$, the more entangled the Z boson pairs is and the larger the value of $I_{3}$ is. This feature is shown in the figure 3. Here the data points represent the expected value of the Bell operator at the specific off-shell Z mass. Clearly, the Bell inequality is violated significantly on the contrary to the predictions of the classical deterministic theories. The actual lower limit of the off-shell Z mass can be larger than 30 GeV. However, putting a lower mass limit on $M_{Z}^{*}$ would increase the statistical uncertainty in the measurement, producing low significance. So we are limited with 30 GeV for the off-shell boson mass.
@@
The final measurements of the observable quantities are shown in  2, 3 and 4 with corresponding CM energies at the four different lower mass limits $M_{Z}^{*}\in[0,10,20,30]$ GeV. The value of $I_{3}$ becomes larger as expected because larger $m_{Z}^{*}$ means the stated entangled more. However, the statistical uncertainties rise as the $M_{Z}^{*}$ mass gets larger. Either non-zero value of the correlation coefficients indicates that the two Z boson states are entangled. Quantum entanglement can be probed up to $4\sigma$ of significance with lower $M_{Z}^{*}$ cut or $2\sigma\sim 3\sigma$ with higher $M_{Z}^{*}$ cut, using either one of the correlation coefficients $C_{2,1,2,-1}$ and $C_{2,2,2,-2}$. The significance of the violation of Bell inequality can be obtained up to $2\sigma$.
@@ IV   Summary and conclusion
In this work we have explored the prospects for probing quantum entanglement and violations of the Bell inequality in $H\to ZZ\to 4\ell$ events at a future muon collider. When the Higgs boson is produced through vector boson fusion, the resulting final state is essentially background-free. Simulations are done at three CM energies, $\sqrt{s}=$ 1, 3, and 10 TeV, and calculations are done for an integrated luminosity of 30 $\rm{ab^{-1}}$. Because the Higgs boson is a spin-zero particle, the $ZZ$ system is in a spin-singlet state, which greatly reduces the number of free parameters in the spin-density matrix. By measuring spin-correlation coefficients, we are able to not only test for the existence of quantum entanglement and but also to test for a much more stringent condition, the violation of the Bell inequality.
The signal events are selected focusing on two electron and two muon channels coming from $ZZ$ decay. Both the on-shell and off-shell $Z$ bosons are identified by determining the invariant mass of these lepton pairs. The Z boson with largest or close to the real mass is tagged as on-shell, while the other one with very small mass compared to the on-shell one is tagged as off-shell. This method is very unique comparing to Z pair production in other processes or colliders. We also included in our simulation the detector effects using Delphes program configured for a muon detector. Cuts are being applied on the lepton kinematic variables $p_{\rm T}$ and $|\eta|$ based on the current muon detection technology. We have also simulated ZZ pair production through $W$ fusion as background to our signal and found out that it to be negligible.
The density matrix of the ZZ state is expressed with a simple yet very effective parametrization [25] which contains only three independent parameters. In the end, we are left with spin-correlation coefficients $C_{2,1,2,-1}$ and $C_{2,2,2,-2}$ that can determine the density matrix. These coefficients can be extracted from actual or simulated data by determining the spherical coordinates of the final leptons, namely electron and muon. Any non-zero value of $C_{2,1,2,-1}$ or $C_{2,2,2,-2}$ signals quantum entanglement between paired Z boson states. Finally, we found that entanglement can be probed with a significance of around $4\sigma$ and the violation of the Bell inequality can be tested up to $2\sigma$ level.

### v2

@@ III   Numerical Simulation and results
We examine the sensitivity of a future Muon Collider to the entanglement and violation of Bell inequality in $H\to ZZ$ process. The corresponding Higgs boson is produced through VBF process: $W^{+}W^{-}\rightarrow H$. It is worth noting that the signal of this VBF process in a Muon Collider is very clean, almost background free as to be shown below. Here the Higgs boson is on-shell and eventually decays into pair of $Z$ bosons , one of which is on-shell while the other is off-shell. We perform a Monte-Carlo simulation of $\mu^{+}\mu^{-}\to\nu_{\mu}\bar{\nu}_{\mu}H$, $H\to ZZ^{*}\to(\ell^{+}\ell^{-})(\ell^{+}\ell^{-})$ events, where $\ell\in\{e,\mu\}$ and $Z^{*}$ representing the off-shell $Z$ boson, using the Madgraph software [39] which includes full spin correlation and Breit-Wigner effects. A sample of one million events are generated at leading order with $\sqrt{s}=1,3,10$ TeV energies for a Muon Collider. These events are then passed to PYTHIA 8.3 [48] for further event hadronization. Eventually all events are passed through Delphes program [40] to include detector effects with a configured file designed for a Muon Collider  [^1] https://github.com/delphes/delphes/blob/master/cards/delphes_card_MuonColliderDet.tcl.
@@
The LO cross sections for Higgs production through VBF is relatively large at a TeV Muon Collider [50], which is about $1{\rm pb}$. However, the final cross sections for the considered process is suppressed due to the smaller branching ratio of the Higgs boson into four leptons (muon or electron pairs). The final leptons coming from the $Z$ boson pairs can be identified as either four electrons, four muons, or two electrons and two muons. We focus on the $e^{+}e^{-}\mu^{+}\mu-$ category for simplicity. We choose not to apply cuts on final leptons to avoid possible bias in quantum entanglement measurement when generating parton-level events with Madgraph . On the other hand, cuts on the kinematic variables such as transverse momentum $p_{\rm T}$ and pseudo-rapidity $\eta_{\ell}$ of leptons are applied during the run of Delphes program to mimic the real experiment detector effect. The minimum value of the electron or muon transverse momentum, $p_{\rm T}$, is set to 0.5 GeV, and the maximum value for absolute rapidity, $|\eta_{\ell}|$, is 2.5.
@@
In the references [7, 26, 42], it is argued that the two $Z$ boson states is not differentiable because of their interaction nature with leptons. However, this difficulty may be avoided in this scenario. The two $Z$ boson can be differentiated. The on-shell $Z$ boson, labeled as $Z_{1}$, can be identified if the invariant mass of lepton pairs is very close to the true physical mass of $Z$ boson. Similarly, the off-shell $Z$ boson, labeled as $Z_{2}$, can also be identified if invariant mass of the lepton pair is much smaller than the actual $Z$ mass.label on-shell $Z$ as $Z_{1}$ whose momentum are determined from summing over its decay products $\ell_{1}^{+}\ell_{1}^{-}$, while the off-shell $Z$ is labeled $Z_{2}$ and its momentum are reconstructed from its decay leptons $\ell_{2}^{+}\ell_{2}^{-}$. We apply identical coordinate system to measure the spin polarization observable as in references [41, 24]. First of all, a set of coordinates are set up in the Higgs boson rest frame into which both boson pairs and leptons are boosted. Finally, the angular coordinates of the leptons, $(\theta_{1}^{-},\phi_{1}^{-})$ for negative-charged lepton from $Z_{1}$ decay and $(\theta_{2}^{-},\phi_{2}^{-})$ for negative-charged lepton from $Z_{2}$ decay can be obtained and used to determine the coefficients for density matrix in Eq. 9. through the integration given in 15.
@@
Figure 3: The expectation value of Bell inequality as function of the off-shell Z mass $M_{Z_{2}}$. $I_{3}$ is measured at three CM energy. Points represent mean value of $I_{3}$.
Backgrounds at muon colliders are generally several orders of magnitude smaller than at hadron colliders [51]. We expect the electroweak process $\mu^{+}\mu^{-}\to ZZ\to 4\ell$ to be the main background in our analysis, where the two Z bosons are produced through W boson fusion. The four-lepton invariant mass distributions for the signal and background are shown in Fig. 2. The red-filled area depicts our signal events while the black-filled area represents the background events. The signal events are clearly distinguishable from the signal events, making it a very good approximation to assume that there is no significant background for our Higgs signal.
The statistical uncertainties of the coefficients $C_{2,1,2,-1},~{}C_{2,2,2,-2}$ and $I_{3}$ are estimated through slicing the total one million events into number of pseudo-experiments, and calculate the observables from the differential distribution. Each pseudo-experiment includes the number of events corresponding to the achievable events for target luminosity and cross section shown in Tab. 1. So that we will have approximately 500 $\sim$ 2000 pseudo-experiments. Repeating the calculation over these pseudo-experiments will give us the mean and standard deviation for each observable. Here, the standard deviation is understand as statistical uncertainty. Because of the clean final states and fine experimental resolutions for light charged leptons, systematic uncertainty are not considered in this study. As stated in previous sections, the entanglement strength is relevant with the off-shell Z boson mass: the larger the mass $M_{Z_{2}}$, the more entangled the $Z$ boson pairs are, and the greater the value of $I_{3}$. This feature is shown in the figure 3. Here the data points represent the mean value of the Bell operator at the specific off-shell Z mass with corresponding CM energies. The blue line depicts the $I_{3}$ value extracted from data with $\sqrt{s}=1$ TeV, while the red and green lines represent $I_{3}$ determined from data with respect 3 and 10 TeV CM energies. Clearly, the Bell inequality is significantly violated, in contrast to to the predictions of the classical deterministic theories. The actual lower limit of the off-shell Z mass can be larger than 30 GeV. However, putting a to much mass limit on $M_{Z_{2}}$ would increase the statistical uncertainty in the measurement, producing low significance. So we are limited with 30 GeV mass-limit for the off-shell boson mass.
@@
The final measurements of the observable quantities are shown in  2, 3 and 4 with corresponding CM energies at the four different lower mass limits $M_{Z_{2}}\in[0,10,20,30]$ GeV. The value of $I_{3}$ becomes larger as expected because larger $M_{Z_{2}}$ means the joint states entangled more. However, the statistical uncertainties rise as the $M_{Z_{2}}$ mass gets larger. Either non-zero value of the correlation coefficients indicates that the two Z boson states are entangled. Quantum entanglement can be probed up to $4\sigma$ of significance with lower $M_{Z_{2}}$ cut or $2\sigma\sim 3\sigma$ with higher $M_{Z_{2}}$ cut, using either one of the correlation coefficients $C_{2,1,2,-1}$ and $C_{2,2,2,-2}$. The significance of the violation of Bell inequality can be obtained up to $2\sigma$.
@@ IV   Summary and conclusion
In this work, we have explored the prospects for probing quantum entanglement and violation of the optimized Bell inequality, CGLMP inequality, in $H\to ZZ\to 4\ell$ process at a future Muon Collider. When the Higgs boson is produced via vector boson fusion, the resulting final state is essentially background-free. Simulations are performed at three center-of-mass energies, $\sqrt{s}$ = 1, 3, and 10 TeV, with assumed integrated luminosity of $30\rm{ab^{-1}}$. Because, the Higgs boson is a spin-zero particle, the $ZZ$ system form a spin-singlet state, significantly reducing the number of free parameters in the spin-density matrix. By measuring spin-correlation coefficients, we are able to not only test for the presence of quantum entanglement but also prove a much more stringent condition: the violation of Bell inequality.
The signal events are selected by focusing on two electron and two muon channels coming from $ZZ$ decay. Both the on-shell and off-shell $Z$ bosons are identified by determining the invariant mass of these lepton pairs. The Z boson with largest or close to the real mass is tagged as on-shell, while the other one, with very small mass, is tagged as off-shell. This method is very unique compared to Z pair production in other processes or colliders. We also included the detector effects in our simulation using Delphes program configured for a muon detector. Cuts are being applied on the lepton kinematic variables $p_{\rm T}$ and $|\eta|$ based on the current muon detection technology. We have also simulated ZZ pair production through $W$ fusion as background to our signal and found out that it to be negligible.
The density matrix of the ZZ state is expressed with a simple yet very effective parametrization [24] which contains only two independent parameters. In the end, we are left with spin-correlation coefficients $C_{2,1,2,-1}$ and $C_{2,2,2,-2}$ that can determine the density matrix. These coefficients can be extracted from actual or simulated data by determining the spherical coordinates of the final leptons, namely electron and muon. Any non-zero value of $C_{2,1,2,-1}$ or $C_{2,2,2,-2}$ signals quantum entanglement between paired Z boson states. Finally, we found that entanglement can be probed with a significance of around $4\sigma$ and the violation of the Bell inequality can be tested up to $2\sigma$ level.